# Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model

### v2

@@ Abstract
Recent studies on topological materials are expanding into the nonlinear regime, while the central principle, namely the bulk-edge correspondence, is yet to be elucidated in the strongly nonlinear regime. Here, we reveal that nonlinear topological edge modes can exhibit the transition to spatial chaos by increasing nonlinearity, which can be a universal mechanism of the breakdown of the bulk-edge correspondence. Specifically, we unveil the underlying dynamical system describing the spatial distribution of zero modes and show the emergence of chaos. We also propose the correspondence between the absolute value of the topological invariant and the dimension of the stable manifold under sufficiently weak nonlinearity. Our results provide a general guiding principle to investigate the nonlinear bulk-edge correspondence that can potentially be extended to arbitrary dimensions.
@@
Introduction.— Band topology is responsible for the existence and absence of zero modes localized at the edge of the sample, which is known as the bulk-edge correspondence Hasan2010 ; Qi2011 . A typical model realizing the nontrivial band topology is the Su-Schrieffer-Heeger (SSH) model Su1979 , which is a one-dimensional tight-binding model with staggered linear couplings. While band topology has been studied mainly in electronic systems such as the celebrated quantum Hall effect Klitzing1980 ; Thouless1982 ; Haldane1988 ; Kane2005 , recent studies have also revealed the existence of topological edge modes in various fields of physics including photonics Haldane2008 ; Khanikaev2013 ; Ozawa2019 , fluid Yang2015 ; Delplace2017 ; Souslov2017 ; Shankar2017 , and cold atoms Atala2013 ; Jotzu2014 . Unlike the standard Schrödinger equation, the dynamics of such classical or quantum bosonic systems are often described by nonlinear equations Gross1961 ; Pitaevskii1961 ; Marchetti2013 ; Buck1988 ; Luo1991 ; Boyd2003 . There are also attempts to extend the notion of topology to such nonlinear systems Lumer2013 ; Leykam2016 ; Harari2018 ; Smirnova2019 ; Zangeneh2019 ; Zhang2020 ; Smirnova2020 ; Ota2020 ; Ivanov2020 ; Lo2021 ; Mukherjee2021 ; Kotwal2021 ; Sone2021 ; Mochizuki2021 ; Li2022 ; Jurgensen2021 ; Mostaan2022 ; Wachtler2023 ; Isobe2023 , which have revealed that nonlinear effects can induce topological phase transitions. Especially, there are several studies on nonlinear SSH models Chen2014 ; Hadad2016 ; Wang2019 ; Xia2020 ; Tuloup2020 ; Ezawa2021 ; Zhou2022 . The emergence of nonlinearity-induced topological edge modes depends on the amplitude Hadad2016 ; Darabi2019 ; Lukas2020 , based on which one can appropriately define the nonlinear topological invariants Tuloup2020 ; Zhou2022 ; Sone2023 . However, in strongly nonlinear regimes, some studies Ezawa2021 ; Fu2022 have pointed out the disappearance of edge modes. Thus, the bulk-edge correspondence in arbitrary strength of nonlinearity remains unelucidated.
@@
In this paper, we reveal that the strong nonlinear effect induces the transition from topological edge modes to spatially chaotic zero modes. Such a chaos transition in zero modes is expected to be a universal mechanism of the breakdown of the bulk-edge correspondence in nonlinear systems. Specifically, we find that the spatial distribution of zero modes is captured by a discrete dynamical system. Focusing on a minimal model of one-dimensional nonlinear topological insulators, we analyze the bifurcation in the corresponding dynamical system and reveal that it shows the period-doubling bifurcation to chaos. The bifurcation point of the period-doubling bifurcation corresponds to the parameter where the bulk-edge correspondence collapses. Concerning the physical meaning of the absolute value of a nonlinear topological invariant, we propose that under sufficiently weak nonlinearity, it corresponds to the dimension of the stable manifold in the dynamical system describing the zero modes. We demonstrate such correspondence in a model with long-range hoppings than the minimal model. Just like the minimal model, the bulk-edge correspondence of higher nonlinear topological invariants can also be broken by the chaos transition. These results can be potentially extended to arbitrary dimensions and thus provide the guiding principle to elucidate the bulk-edge correspondence and its breakdown in nonlinear topological insulators.
@@
Figure 1: Schematics of the long-range nonlinear SSH model. The model has two sublattices (red and blue circles) and staggered nearest hopping. The strength of the intercell nearest-neighbor hopping depends on the state variables. We also introduce long-range hoppings (green curves) to investigate the effect of higher winding numbers.
@@
Setup.— Following some previous studies Bomantara2017 ; Tuloup2020 ; Zhou2022 ; Sone2023 ; Li2022 , we define the nonlinear eigenvalue problem to extend the topological invariants and edge modes to nonlinear lattice systems. Labeling the sites and internal degrees of freedom by $x$ and $j$, we consider the general dynamics $i\partial_{t}\Psi_{j}(x)=f_{j}(\bm{\Psi},x)$ with $f_{j}$ ($j=1,\cdots,M$) being nonlinear functions. Here $\bm{\Psi}$ is the abbreviation for $\{\Psi_{j}(x)\}_{x,j}$. We also assume the $U(1)$ symmetry, $f_{j}(e^{i\theta}\bm{\Psi})=e^{i\theta}f_{j}(\bm{\Psi})$ ($j=1,\cdots,M$), and the translation invariance, which also exist in prototypical setups of linear topological insulators Thouless1982 ; Haldane1988 ; Kane2005 . By imposing these symmetries, we can clearly identify the wavenumber-space description of the nonlinear system as below.
@@
Corresponding to the nonlinear dynamics, we formulate the nonlinear eigenvalue problem by defining a nonlinear eigenvector $\bm{\Psi}$ and a nonlinear eigenvalue $E$ as a vector and a scalar satisfying $f_{j}(\bm{\Psi},x)=E\Psi_{j}(x)$. We note that the nonlinear eigenvector $\bm{\Psi}$ corresponds to a periodically oscillating steady state $\Psi_{j}(x;t)=e^{-iEt}\Psi_{j}(x)$ in the original nonlinear dynamics. Furthermore, assuming the Bloch ansatz $\bm{\Psi}(x)=e^{ikx}\bm{\psi}(k)$, one can derive the wavenumber-space description of the nonlinear eigenequation
@@
Finally, we define the nonlinear topological invariants by substituting linear eigenvectors with nonlinear ones in the definitions of conventional topological invariants.
We next explicitly define the nonlinear winding number characterizing a one-dimensional nonlinear topological insulator with the sublattice symmetry Altland1997 ; Kitaev2009 ; Ryu2010 (see Methods) and its bulk-edge correspondence. We assume that the nonlinear eigenequation (Eq. (1)) has the form of
@@
where $q(\bm{\psi},k)$ is a matrix parametrized by the wavenumber $k$ and the nonlinear eigenvector $\bm{\psi}=(\bm{\psi}_{A},\bm{\psi}_{B})$. In this manuscript, we further focus on the case that $q(\bm{\psi},k)$ only depends on the wavenumber and the amplitude of the state $||\bm{\psi}_{A}||^{2}+||\bm{\psi}_{B}||^{2}$, $q(\bm{\psi},k)=q(||\bm{\psi}_{A}||^{2}+||\bm{\psi}_{B}||^{2},k)$ with |||||$\cdot$||| | ⋅ | | being the vector norm. As in a previous study Sone2023 , we consider special solutions where the amplitude is fixed $||\bm{\psi}_{A}||^{2}+||\bm{\psi}_{B}||^{2}=w$ independently of $k$. Then, the nonlinear winding number is defined as
@@
The amplitude dependence of the nonlinear winding number implies the potential nonlinearity-induced topological phase transition.
@@
Regarding the bulk-edge correspondence of this nonlinear winding number, we find that the nonzero (zero) winding number basically corresponds to the existence (absence) of the localized zero modes when we identify the amplitude $w$ in Eq. (9) to be the edge amplitude, $\sum_{j}|\Psi_{j}(1)|^{2}=w$. Unlike linear systems, one can obtain zero modes even in trivial phases, while they are anti-localized. More specifically, in semi-infinite systems, we fix the edge amplitude $||\Psi(x=1)||^{2}$ of the zero mode to be $w$. Then, the zero mode should exhibit localization (resp. delocalization) in the case of $\nu(w)\neq 0$ (resp. $\nu(w)=0$). However, such a bulk-edge correspondence can be broken by the transition to chaos as we discuss later.
@@
Before moving to the breakdown of the bulk-edge correspondence, we discuss the triviality of the anti-localized zero modes in more detail. First, we can define a topologically trivial phase in nonlinear systems as a phase that can be adiabatically deformed into a linear trivial phase. We can show the existence of such an adiabatic deformation of the phase with the anti-localized zero mode into a linear trivial phase from the fact that one can deform the anti-localized zero mode into a diverging mode by continuously eliminating the nonlinear terms. Since such diverging modes are not regarded as physically feasible edge modes, the divergence indicates the absence of zero modes after adiabatically deforming the nonlinear system into a linear system. Therefore, the anti-localized zero modes are regarded as topologically trivial modes.
@@
Figure 2: Bifurcation diagram of zero modes in the nonlinear SSH model. The blue curves and dots represent the absolute values of zero modes at each lattice site far from the open boundary. The red dashed lines separate the regions of weakly nonlinear topological phase, nonlinearity-induced topological phase, periodic zero mode phase, and chaotic zero mode phase. The light blue area corresponds to the parameter region where the nonlinear winding number becomes one. The upper bound of the light blue area (the red curves) corresponds to the unstable fixed points. The green curve shows the lower bound of the light blue area at $a>2$, while it does not correspond to the convergent value of the dynamical system, which indicates the breakdown of the bulk-boundary correspondence. The bulk-edge correspondence holds only at $a<2$, where steady solutions are obtained. The parameters used are $b=c=-1$, $d=1$. We set the initial condition $\Psi_{A}(0)=0.1$.
@@
Figure 3: Cobweb plots and spatial distributions of zero modes in the nonlinear SSH model. a-d Cobweb plots at different $a$ are shown. The blue curves represent the absolute values of the nonlinear functions $F(\Psi)$ in Eq. (19). The green lines represent $|F(\Psi)|=\Psi$. The red lines show the dynamics of Eq. (19), which correspond to the blue polylines. a Weakly nonlinear topological case. If the nonlinear winding number is $\nu=1$ in the linear limit, localized edge modes converging to zero are obtained for smaller initial amplitudes than that represented by the red dotted line. We use the parameter $a=0.8$. b Nonlinearity-induced topological phase. When the nonlinearity-induced topological phase transition from a trivial phase to a topological phase occurs, localized zero modes are obtained if the initial amplitude is in the region sandwiched by the red dotted lines. We use the parameter $a=1.8$. c Periodic zero mode. We consider $a=2.1$ and obtain a stable periodic orbit. d Chaotic zero mode. At large $a$ ($a=2.5$ in this panel), we obtain a chaotic dynamics of a zero mode. e-h Zero modes at each parameter are shown. The colored polylines show the spatial distributions of zero modes, where the color shows the difference in the initial condition $\Psi_{A}(1)$. The red dashed lines show the critical amplitude of the nonlinear winding number. We use the parameters $a=0.8$, $1.8$, $2.1$, $2.5$ in panels e, f, g, h for each, which correspond to the upper panels a-d.
Nonlinear SSH model.— To investigate the nonlinear effects on topological edge modes, we consider the nonlinear SSH model, which is a minimal model of one-dimensional nonlinear topological insulators. The nonlinear SSH model has two sublattices labeled A and B (cf. Fig. 1), and its dynamics is described as
@@
where $\Psi_{A(B)}(x)$ represents the state variables at the A (B) sublattice of the $x$th unit cell. In the following, we focus on the case that $a$, $b$, $c$, and $d$ are real, and $b$ is equal to $c$.
The wavenumber-space description of the nonlinear SSH model has the form of Eq. (8):
@@
Figure 4: Vector fields representing the deviation of state variables in the extended nonlinear SSH model. a Phase diagram of the linear extended SSH model. Each color represents the parameter regions with the same winding number, which is shown by the numbers in the regions. b-d Vector field and the stable manifold. The blue curved arrows represent the vector field. The red filled circles are stable fixed points or saddle points, and the green circle is a fully unstable fixed point. The green curves are eye-guides of the one-dimensional stable manifolds. b If the winding number is one under weak nonlinearity (the blue circle in panel a), the stable manifold is one-dimensional. The parameters used are $a=0.5$, $b=d=-1$, and $\alpha=0.25$. c When the winding number is two under weak nonlinearity (the red square in panel a), the fixed point is stable and thus has a two-dimensional stable manifold. The parameters used are $a=1$, $b=d=-1$, and $\alpha=4$. d When the nonlinearity-induced topological phase transition from zero to one occurs (the green arrow in panel a), nonzero fixed points appear and their stable manifold is one-dimensional. The parameters used are $a=1.75$, $b=d=-1$, and $\alpha=0.25$.
Bifurcation and spatially chaotic zero modes.— We analyze the zero mode of the nonlinear SSH model and reveal the breakdown of the bulk-edge correspondence between edge modes and the nonlinear winding number. Specifically, we consider the right semi-infinite system which has an open boundary at $x=1$. Then, the zero mode has zero amplitude on the B sublattice, i.e., $\Psi_{B}(x)=0$, and the spatial distribution on the A sublattice is described by the following nonlinear dynamical system:
@@
where $F(\Psi_{A}(x))$ is the nonlinear function determining the spatial distribution and independent of $\Psi_{B}(x)$ and $c$. As discussed above, the bulk-edge correspondence implies that the nonzero winding number corresponds to the existence of the localized zero modes. On the other hand, the nonlinear dynamical system (Eq. (19)) is known as the cubic map Rogers1983 ; Skjolding1983 and shows bifurcations to chaos. We find that this bifurcation leads to the breakdown of the bulk-edge correspondence.
We numerically demonstrate the bifurcation in the zero mode of the nonlinear SSH model by using a bifurcation diagram, which represents the behavior of $\Psi_{A}(x)$ far from the edge $x\rightarrow\infty$. In more detail, the bifurcation diagram shows the values of $\Psi_{A}(x)$ at large $x$ for different $a$’s. If we obtain only one value at a fixed $a$ in the bifurcation diagram, the dynamical system (Eq. (19)) converges to that value in the limit of $x\rightarrow\infty$. Before a chaos transition, the convergence of the dynamical system is destabilized at a critical value of $a$, and we obtain stable periodic solutions. Such periodic solutions are represented by two or more curves in the bifurcation diagram. After a chaos transition, one obtains scattered points at a fixed $a$, which indicates the unstable and unpredictable dynamics of chaos.
Figure 2 shows the bifurcation diagram of Eq. (19) at $b=-1$, $d=1$ (see Methods for the numerical method). At $a<2$, we confirm the convergence to a steady value. In particular, in the case of $21<a<$ 2, the convergent value corresponds to the critical amplitude where the nonlinear winding number is changed. This correspondence is not a coincidence, because the boundary between topological and trivial phases must be a threshold where the localization or delocalization of a zero mode is switched, and thus a spatially uniform mode must appear at the boundary, which is a steady state of Eq. (19). In more detail, when we fix the absolute value at the boundary as $|\Psi_{A}(1)|^{2}=w$, we obtain the following equation from Eq. (19):

### v3

@@ Abstract
Recent studies on topological materials are expanding into the nonlinear regime, while the central principle, namely the bulk–edge correspondence, is yet to be elucidated in the strongly nonlinear regime. Here, we reveal that nonlinear topological edge modes can exhibit the transition to spatial chaos by increasing nonlinearity, which can be a universal mechanism of the breakdown of the bulk–edge correspondence. Specifically, we unveil the underlying dynamical system describing the spatial distribution of zero modes and show the emergence of chaos. We also propose the correspondence between the absolute value of the topological invariant and the dimension of the stable manifold under sufficiently weak nonlinearity. Our results provide a general guiding principle to investigate the nonlinear bulk–edge correspondence that can potentially be extended to arbitrary dimensions.
@@
Introduction.— Band topology is responsible for the existence and absence of zero modes localized at the edge of the sample, which is known as the bulk–edge correspondence Hasan2010 ; Qi2011 . A typical model realizing the nontrivial band topology is the Su–Schrieffer–Heeger (SSH) model Su1979 , which is a one–dimensional tight–binding model with staggered linear couplings. While band topology has been studied mainly in electronic systems such as the celebrated quantum Hall effect Klitzing1980 ; Thouless1982 ; Haldane1988 , recent studies have also revealed the existence of topological edge modes in various fields of physics including photonics Haldane2008 ; Khanikaev2013 , fluid Yang2015 , and cold atoms Atala2013 . Unlike the standard Schrödinger equation, the dynamics of such classical or quantum bosonic systems are often described by nonlinear equations Gross1961 ; Pitaevskii1961 ; Marchetti2013 ; Luo1991 ; Boyd2003 . There are also attempts to extend the notion of topology to such nonlinear systems Lumer2013 ; Leykam2016 ; Harari2018 ; Smirnova2019 ; Zangeneh2019 ; Zhang2020 ; Smirnova2020 ; Ota2020 ; Ivanov2020 ; Lo2021 ; Mukherjee2021 ; Kotwal2021 ; Sone2021 ; Mochizuki2021 ; Li2022 ; Jurgensen2021 ; Mostaan2022 ; Wachtler2023 ; Isobe2023 , which have revealed that nonlinear effects can induce topological phase transitions. Especially, there are several studies on nonlinear SSH models Chen2014 ; Hadad2016 ; Wang2019 ; Xia2020 ; Tuloup2020 ; Ezawa2021 ; Zhou2022 ; Engelhardt2017 ; Chaunsali2021 ; Ma2021 ; Jezequel2022 ; Ma2023 . The emergence of nonlinearity–induced topological edge modes depends on the amplitude Hadad2016 ; Darabi2019 ; Lukas2020 , based on which one can appropriately define the nonlinear topological invariants Tuloup2020 ; Zhou2022 ; Sone2023 . However, in strongly nonlinear regimes, some studies Ezawa2021 ; Fu2022 have pointed out the disappearance of edge modes. Thus, the bulk–edge correspondence in arbitrary strength of nonlinearity remains unelucidated.
@@
In this paper, we reveal that the strong nonlinear effect induces the transition from topological edge modes to spatially chaotic zero modes. Such a chaos transition in zero modes is expected to be a universal mechanism of the breakdown of the bulk–edge correspondence in nonlinear systems. Specifically, we find that the spatial distribution of zero modes is captured by a discrete dynamical system. Focusing on a minimal model of one–dimensional nonlinear topological insulators, we analyze the bifurcation in the corresponding dynamical system and reveal that it shows the period–doubling bifurcation to chaos. The bifurcation point of the period–doubling bifurcation corresponds to the parameter where the bulk–edge correspondence collapses. Concerning the physical meaning of the absolute value of a nonlinear topological invariant, we propose that under sufficiently weak nonlinearity, it corresponds to the dimension of the stable manifold in the dynamical system describing the zero modes. We demonstrate such correspondence in a model with long–range hoppings than the minimal model. Just like the minimal model, the bulk–edge correspondence of higher nonlinear topological invariants can also be broken by the chaos transition. These results can be potentially extended to arbitrary dimensions and thus provide the guiding principle to elucidate the bulk–edge correspondence and its breakdown in nonlinear topological insulators.
@@
Figure 1: Schematics of the long-range nonlinear SSH model. The model has two sublattices (red and blue circles) and staggered nearest hopping. The strength of the intercell nearest–neighbor hopping depends on the state variables as $a+b|\Psi_{A}(x)|^{2}+c|\Psi_{B}(x)|^{2}$ with $a$, $b$, and $c$ being real parameters. $d$ is also a real parameter that determines the strength of intracell nearest–neighbor hopping. We also introduce long–range hoppings (green curves, the strength $\alpha$) to investigate the effect of higher winding numbers.
@@
Setup.— Following some previous studies Bomantara2017 ; Tuloup2020 ; Zhou2022 ; Sone2023 ; Li2022 , we define the nonlinear eigenvalue problem to extend the topological invariants and edge modes to nonlinear lattice systems. Labeling the sites and internal degrees of freedom by $x$ and $j$, we consider the general dynamics $i\partial_{t}\Psi_{j}(x)=f_{j}(\bm{\Psi},x)$ with $f_{j}$ ($j=1,\cdots,M$) being nonlinear functions. Here $\bm{\Psi}$ is the abbreviation for $\{\Psi_{j}(x)\}_{x,j}$. We also assume the $U(1)$ symmetry, $f_{j}(e^{i\theta}\bm{\Psi})=e^{i\theta}f_{j}(\bm{\Psi})$ ($j=1,\cdots,M$), and the translation invariance, which also exist in prototypical setups of linear topological insulators Thouless1982 ; Haldane1988 . By imposing these symmetries, we can clearly identify the wavenumber–space description of the nonlinear system as below.
@@
Corresponding to the nonlinear dynamics, we formulate the nonlinear eigenvalue problem by defining a nonlinear eigenvector $\bm{\Psi}$ and a nonlinear eigenvalue $E$ as a vector and a scalar satisfying $f_{j}(\bm{\Psi},x)=E\Psi_{j}(x)$. We note that the nonlinear eigenvector $\bm{\Psi}$ corresponds to a periodically oscillating steady state $\Psi_{j}(x;t)=e^{-iEt}\Psi_{j}(x)$ in the original nonlinear dynamics. Furthermore, assuming the Bloch ansatz $\bm{\Psi}(x)=e^{ikx}\bm{\psi}(k)$, one can derive the wavenumber–space description of the nonlinear eigenequation
@@
Finally, we define the nonlinear topological invariants by substituting linear eigenvectors with nonlinear ones in the definitions of conventional topological invariants. We note that a previous paper Isobe2023 discussed the bulk–edge correspondence in nonlinear eigenvalue problems with respect to eigenfrequencies, which however describe linear dynamics of higher–order differential equations. In contrast, we here focus on the situation in which the dynamics itself is nonlinear.
We next explicitly define the nonlinear winding number characterizing a one–dimensional nonlinear topological insulator with the sublattice symmetry Altland1997 ; Kitaev2009 ; Ryu2010 ; Jezequel2022 (see Methods) and its bulk-edge correspondence. We assume that the nonlinear eigenequation (Eq. (1)) has the form of
@@
where $q(\bm{\psi},k)$ is a matrix parametrized by the wavenumber $k$ and the nonlinear eigenvector $\bm{\psi}=(\bm{\psi}_{A},\bm{\psi}_{B})$. In this manuscript, we further focus on the case that $q(\bm{\psi},k)$ only depends on the wavenumber and the amplitude of the state $||\bm{\psi}_{A}||^{2}+||\bm{\psi}_{B}||^{2}$, $q(\bm{\psi},k)=q(||\bm{\psi}_{A}||^{2}+||\bm{\psi}_{B}||^{2},k)$ with |||||$\cdot$||| | ⋅ | | being the vector norm. As in a previous study Sone2023 , we consider special solutions where the amplitude is fixed $||\bm{\psi}_{A}||^{2}+||\bm{\psi}_{B}||^{2}=w$ independently of $k$. This assumption is natural because the amplitude is a conserved quantity that one can control by tuning the energy injected to excite the initial state. Then, the nonlinear winding number is defined as
@@
The amplitude dependence of the nonlinear winding number implies the potential nonlinearity–induced topological phase transition.
@@
Regarding the bulk–edge correspondence of this nonlinear winding number, we find that the nonzero (zero) winding number basically corresponds to the existence (absence) of the localized zero modes (i.e., nonlinear eigenvectors with zero eigenvalue localized at the edge) when we identify the amplitude $w$ in Eq. (9) to be the edge amplitude, $\sum_{j}|\Psi_{j}(1)|^{2}=w$. We here assume the zero eigenvalue of topological edge modes because of the sublattice symmetry of the system. Unlike linear systems, one can obtain zero modes even in trivial phases, while they are anti–localized. More specifically, in semi–infinite systems, we fix the edge amplitude $||\Psi(x=1)||^{2}$ of the zero mode to be $w$. Then, the zero mode should exhibit localization (resp. anti–localization) in the case of $\nu(w)\neq 0$ (resp. $\nu(w)=0$). However, such a bulk–edge correspondence can be broken by the transition to chaos as we discuss later.
@@
Before moving to the breakdown of the bulk–edge correspondence, we discuss the triviality of the anti–localized zero modes in more detail. First, we can define a topologically trivial phase in nonlinear systems as a phase that can be adiabatically deformed into a linear trivial phase. We can show the existence of such an adiabatic deformation of the phase with the anti–localized zero mode into a linear trivial phase from the fact that one can deform the anti–localized zero mode into a diverging mode by continuously eliminating the nonlinear terms. Since such diverging modes are not regarded as physically feasible edge modes, the divergence indicates the absence of zero modes after adiabatically deforming the nonlinear system into a linear system. Therefore, the anti–localized zero modes are regarded as topologically trivial modes.
@@
Figure 2: Bifurcation diagram of zero modes in the nonlinear SSH model. The horizontal axis represents the parameter $a$ determining the strength of nonlinearity. The blue curves and dots represent the absolute values of zero modes $|\Psi|$ at each lattice site far from the open boundary. The red dashed lines separate the regions of weakly nonlinear topological phase, nonlinearity–induced topological phase, periodic zero mode phase, and chaotic zero mode phase. As shown in a previous paper Sone2023 for two–dimensional systems, the bulk–edge correspondence is valid from the weak to the mildly strong nonlinear regimes, by introducing the nonlinear extension of the topological number. However, in the truly strong nonlinear regime, the bulk–edge correspondence is broken down by the bifurcation and the subsequent chaos transition. The boundaries at $a=1,2$ are analytically obtained from the linear stability analysis of the spatial dynamics, and the boundary around $a=2.31$ is numerically estimated from the Lyapunov exponent. The light blue area corresponds to the parameter region where the nonlinear winding number $\nu$ becomes one, whose boundaries are analytically obtained. The upper bound of the light blue area (the red curves) corresponds to the unstable fixed points at $|\Psi_{A}(x)|=\sqrt{a+1}$. The orange curve shows the lower bound of the light blue area ($|\Psi_{A}(x)|=\sqrt{a-1}$) at $a>2$, while it does not correspond to the convergent value of the dynamical system, which indicates the breakdown of the bulk-boundary correspondence. The bulk–edge correspondence holds only at $a<2$, where steady solutions are obtained. The parameters used are $b=c=-1$, $d=1$. We set the initial condition $\Psi_{A}(0)=0.1$.
@@
Figure 3: Cobweb plots and spatial distributions of zero modes in the nonlinear SSH model. a–d Cobweb plots at different strengths of nonlinearity $a$ are shown. The blue curves represent the absolute values of the nonlinear functions $F(\Psi)$ in Eq. (19) with $\Psi$ corresponding to a component of an eigenvector. The orange lines represent $|F(\Psi)|=\Psi$. The red lines show the dynamics of Eq. (19), which correspond to the blue polylines. a Weakly nonlinear topological case. If the nonlinear winding number is $\nu=1$ in the linear limit, localized edge modes converging to zero are obtained for smaller initial amplitudes than that represented by the red dotted line. We use the parameter $a=0.8$. b Nonlinearity–induced topological phase. When the nonlinearity-induced topological phase transition from a trivial phase to a topological phase occurs, localized zero modes are obtained if the initial amplitude is in the region sandwiched by the red dotted lines. We use the parameter $a=1.8$. c Periodic zero mode. We consider $a=2.1$ and obtain a stable periodic orbit. d Chaotic zero mode. At large $a$ ($a=2.5$ in this panel), we obtain a chaotic dynamics of a zero mode. e–h Zero modes at each parameter are shown. The colored polylines show the spatial distributions of zero modes, where the color shows the difference in the initial condition $\Psi_{A}(1)$. The red dashed lines show the critical amplitude of the nonlinear winding number. We use the parameters $a=0.8$, $1.8$, $2.1$, $2.5$ in panels e, f, g, h for each, which correspond to the upper panels a–d.
Nonlinear SSH model.— To investigate the nonlinear effects on topological edge modes, we consider a nonlinear SSH model, which is a minimal model of one–dimensional nonlinear topological insulators. While previous studies have proposed various types of nonlinear extensions of the SSH model Chen2014 ; Hadad2016 ; Wang2019 ; Xia2020 ; Tuloup2020 ; Ezawa2021 ; Zhou2022 ; Engelhardt2017 ; Chaunsali2021 ; Ma2021 ; Jezequel2022 ; Ma2023 , we consider one of its variants that is suitable to demonstrate the chaos transition. Specifically, the nonlinear SSH model considered here has two sublattices labeled A and B (cf. Fig. 1), and its dynamics is described as
@@
where $\Psi_{A(B)}(x)$ represents the state variables at the A (B) sublattice of the $x$th unit cell. In the following, we focus on the case that $a$, $b$, $c$, and $d$ are real, and $b$ is equal to $c$. We here consider an off–diagonal nonlinearity because it preserves the sublattice symmetry and realizes the nonlinearity–induced topological phase transition Hadad2016 ; Sone2023 . One can experimentally realize such an off–diagonal nonlinearity by utilizing, e.g., phase shifts by the Kerr nonlinearity in optical fibers Bisianov2019 or the nonlinear circuit elements Zhou2022 (cf. Supplementary Note 1).
To calculate the nonlinear topological invariant (Eq. (9)), we derive the wavenumber–space description of the nonlinear eigenvalue problem by assuming the Bloch ansatz. The wavenumber–space description of the nonlinear SSH model has the form of Eq. (8):
@@
Figure 4: Vector fields representing the deviation of state variables in the extended nonlinear SSH model. a Phase diagram of the linear extended SSH model. $a$, $d$, and $\alpha$ are real parameters that determine the strength of intercell, nearest-neighbor intracell, and long-range hoppings, respectively. Each color represents the parameter regions with the same winding number, which is shown by the numbers in the regions. b–d Vector field and the stable manifold. The blue curved arrows represent the vector field at $(\Psi(x),\Psi(x+1))$ corresponding to the values of eigenvectors at the sites $x$ and $x+1$. The red disks are stable fixed points, the red squares are saddle points, and the green circle is a fully unstable fixed point. The orange curves are eye–guides of the one–dimensional stable manifolds. b If the winding number is one under weak nonlinearity (the blue circle in panel a), the stable manifold is one-dimensional. The parameters used are $a=0.5$, $b=d=-1$, and $\alpha=0.25$. c When the winding number is two under weak nonlinearity (the red filled square in panel a), the fixed point is stable and thus has a two–dimensional stable manifold. The parameters used are $a=1$, $b=d=-1$, and $\alpha=4$. d When the nonlinearity–induced topological phase transition from zero to one occurs (the green arrow in panel a), nonzero fixed points appear and their stable manifold is one–dimensional. The parameters used are $a=1.75$, $b=d=-1$, and $\alpha=0.25$.
Bifurcation and spatially chaotic zero modes.— We analyze the zero mode of the nonlinear SSH model and reveal the breakdown of the bulk–edge correspondence between edge modes and the nonlinear winding number. Specifically, we consider the right semi-infinite system which has an open boundary at $x=1$. Then, the zero mode has zero amplitude on the B sublattice, i.e., $\Psi_{B}(x)=0$, and the spatial distribution on the A sublattice is described by the following nonlinear dynamical system:
@@
where $F(\Psi_{A}(x))$ is the nonlinear function determining the spatial distribution and independent of $\Psi_{B}(x)$ and $c$. As discussed above, the bulk–edge correspondence implies that the nonzero winding number corresponds to the existence of the localized zero modes, which is defined as the nonlinear eigenvector with zero eigenvalue and exhibiting a larger amplitude at the edge than in the bulk. We here calculate the zero mode from Eq. (19) and confirm its localization by comparing the edge amplitude with the bulk amplitude; we judge that the zero mode is localized if $|\Psi_{A}(1)|>\lim_{x\rightarrow\infty}|\Psi_{A}(x)|$. On the other hand, the nonlinear dynamical system (Eq. (19)) is known as the cubic map Rogers1983 ; Skjolding1983 (see Methods) and shows bifurcations to chaos. We find that this bifurcation leads to the breakdown of the bulk–edge correspondence.
We numerically demonstrate the bifurcation in the zero mode of the nonlinear SSH model by using a bifurcation diagram, which represents the behavior of the dynamical system in Eq. (19) after the relaxation to a steady, periodic, or chaotic solution. In more detail, the bifurcation diagram shows the values of $\Psi_{A}(x)$ at large $x$ for different $a$’s (we consider $100009900<x<$ 10000 and $30<a<$ 3 in Fig. 2). If we obtain only one value at a fixed $a$ in the bifurcation diagram, the dynamical system (Eq. (19)) converges to that value in the limit of $x\rightarrow\infty$. Before a chaos transition, the convergence of the dynamical system is destabilized at a critical value of $a$, and we obtain stable periodic solutions. Such periodic solutions are represented by two or more curves in the bifurcation diagram. After a chaos transition, one obtains scattered points at a fixed $a$, which indicates the unstable and unpredictable dynamics of chaos.
Figure 2 shows the bifurcation diagram of Eq. (19) at $b=-1$, $d=1$ (see Methods for the numerical method). At $a<2$, we confirm the convergence to a steady value. In particular, in the case of $21<a<$ 2, the convergent value corresponds to the critical amplitude where the nonlinear winding number is changed. This correspondence is not a coincidence, because the boundary between topological and trivial phases must be a threshold where the localization or anti–localization of a zero mode is switched, and thus a spatially uniform mode must appear at the boundary, which is a steady state of Eq. (19). In more detail, when we fix the absolute value at the boundary as $|\Psi_{A}(1)|^{2}=w$, we obtain the following equation from Eq. (19):